OpenStudy (anonymous):

FIve coins are tossed. What is the probability that the first two will be heads?

OpenStudy (anonymous):

The combination of 5 coins is 32, because for each coin there are 2 choice, heads or tails. 5 coins=2*2*2*2*2=32 The chance of 2 coins falling heads leaves the choice of 3 coins, which is 2*2*2=8 The probabilty is 8/32=1/4=0.25